### v1

# Is Mamba Effective for Time Series Forecasting?

Abstract
In the realm of time series forecasting (TSF), the Transformer has consistently demonstrated robust performance due to its ability to focus on the global context and effectively capture long-range dependencies within time, as well as discern correlations between multiple variables. However, due to the inefficiencies of the Transformer model and questions surrounding its ability to capture dependencies, ongoing efforts to refine the Transformer architecture persist. Recently, state space models (SSMs), e.g. Mamba, have gained traction due to their ability to capture complex dependencies in sequences, similar to the Transformer, while maintaining near-linear complexity. In text and image tasks, Mamba-based models can improve performance and cost savings, creating a win-win situation. This has piqued our interest in exploring SSM’s potential in TSF tasks. In this paper, we introduce two straightforward SSM-based models for TSF, S-Mamba and D-Mamba, both employing the Mamba Block to extract variate correlations. Remarkably, S-Mamba and D-Mamba achieve superior performance while saving GPU memory and training time. Furthermore, we conduct extensive experiments to delve deeper into the potential of Mamba compared to the Transformer in the TSF, aiming to explore a new research direction for this field. Our code is available at https://github.com/wzhwzhwzh0921/S-D-Mamba.
Introduction
Time series forecasting (TSF) involves leveraging information from past events to predict conditions in the future. These events often have built-in patterns associated with time or variates, for example, the morning and evening peak patterns in traffic issues, and the pattern between temperature and humidity in weather predictions. This naturally leads to the consideration of identifying long-range dependencies (LRD) or variates correlations (VC) from these data to effectively identify and model these patterns for better forecasting. Transformer (Vaswani et al. 2017), which shines brightly in natural language processing (Wolf et al. 2019, 2020) and visual tasks (Liu et al. 2021; Arnab et al. 2021), is based on the self-attention structure, which can encompass the context and extract deep connections between time points in a sequence. Therefore, it is also applied in TSF tasks. Numerous Transformer-based models with impressive capabilities have been introduced (Wu et al. 2021; Zhou et al. 2022), yet the Transformer architecture faces distinct challenges. Foremost is its quadratic computational complexity, leading to a dramatic increase in calculations for long input sequences. Numerous models have attempted to reduce computational complexity by modifying the Transformer’s structure, such as focusing only on a portion of the sequence (Kitaev, Kaiser, and Levskaya 2020; Zhou et al. 2021; Li et al. 2019). However, the loss of node information in this process may also lead to certain performance degradations. A more promising approach involves using linear models (Li et al. 2023a; Zeng et al. 2023), which possesses linear computational complexity. However, compared with the state-of-the-art Transformer models, these models do not perform so well due to their inability to capture LRD or VC. Additionally, since they do not predict trends and rely solely on numerical calculations, accurate predictions can only be achieved when sufficient input information is available (Zeng et al. 2023).
It seems that capturing browsing context, building LRDs or VCs and linear complexity is not achievable simultaneously. The State Space Models (SSM) (Gu, Goel, and Ré 2021; Smith, Warrington, and Linderman 2022) demonstrate potential in simultaneously optimizing both aspects. It employs convolutional calculation to capture LRDs and eliminates hidden states making it benefit from parallel computing and achieving approximately linear complexity in processing speed. Consequently, SSM is inherently well-suited for TSF tasks. A Previous study (Rangapuram et al. 2018) also attempted to employ SSM for TSF, but the architecture is unable to identify and filter content effectively, and the dependencies it captures are solely based on distance, resulting in suboptimal performance. Recent work, Mamba (Gu and Dao 2023), introduces a selective mechanism into SSM, enabling it to discern the importance of information like the attention mechanism. This development has prompted researchers to build Mamba-based models in domains where Transformer has demonstrated superior performance. The outcomes indicate that Mamba exhibits significant potential in both text and image fields, frequently achieving a win-win situation in terms of model performance and computational efficiency (Zhu et al. 2024; Yang, Xing, and Zhu 2024). This improvement is attributed to Mamba’s enhanced capability to process long sequence data, making it inherently suitable for TSF tasks and sparking our interest in exploring Mamba’s application in this area.
As a result, we launch Simple-Mamba (S-Mamba) and Dual-Mamba (D-Mamba) in an attempt to reconcile the performance and speed disputes in the TSF field and open up new research directions for it. S-Mamba employs one Mamba block to process VC, while D-Mamba incorporates an additional mamba block compared to S-Mamba for VC. The distinction between the additional Mamba block in D-Mamba and the original Mamba block lies in the former’s sensitivity to variable correlations (VC) is higher by adjusting the parameters of the Mamba block. In our design, this modification serves two purposes. Firstly, it is akin to multi-head attention mechanism, aiming to capture multiple facets of VC. Secondly, it is intended to enable the model to learn higher-level VC based on the two different sensitivities. Ultimately, our experimental results demonstrate that S-Mamba and D-Mamba not only achieve significant savings in GPU memory usage and training time but also maintain superior performance compared to state-of-the-art models. our contributions can be categorized into the following three aspects:
  • We introduce S-Mamba and D-Mamba, utilizing the SSM mechanism to capture variate correlations. S-Mamba only has one Mamba block, while D-Mamba uses two Mamba blocks with different sensitivity.
  • We evaluate of S-Mamba and D-Mamba. They not only reduce the time and GPU memory required for forecasts but also maintain superior performance.
  • We conduct extensive experiments to further delve deeper into Mamba’s potential in TSF tasks, aiming to explore a new research direction in this field.
Related Work
In conjunction with our work, two main areas of related work were investigated: (1) time series forecasting, and (2) applications of Mamba.
Time Series Forecasting
Recently, there have been two main architectures for TSF approaches, which are Transformer-based and multilayer perceptrons (MLP)-based.
Transformers are primarily designed for tasks that involve processing and generating sequences of tokens (Vaswani et al. 2017). The excellent performance of Transformer-based models has also attracted numerous researchers to focus on time series forecasting tasks (Ahmed et al. 2023). The transformer is utilized by Duong-Trung, Nguyen, and Le-Phuoc (2023) to solve the persistent challenge of long multi-horizon time series forecasting. Time Absolute Position Encoding (tAPE) and Efficient implementation of Relative Position Encoding (eRPE) are proposed in (Foumani et al. 2024) to solve the position encoding problem encountered by Transformer in multivariate time series classification (MTSC). Some researchers have also considered the application of Transformer-based time series forecasting models in specific domains, such as piezometric level prediction (Mellouli, Rabah, and Farah 2022), forecasting crude oil returns (Abdollah Pour, Hajizadeh, and Farineya 2022), predicting the power generation by solar panels (Sherozbek et al. 2023), etc.
While they excel at capturing long-range dependencies in text, they may not be as effective in modeling sequential patterns. The use of content-based attention in Transformers is not effective in detecting essential temporal dependencies, especially for time-series data with weakening dependencies over time and strong seasonality patterns (Woo et al. 2022). Particularly, the predictive capability and robustness of Transformer-based models may decrease rapidly when the input sequence is too long (Wen et al. 2023). Moreover, the $O(\text{N}^{2})$ time complexity makes Transformer-based models cost more computation and GPU memory resources. In addition, the previously mentioned issue of position encoding is also a challenge that deserves attention.
In addition to Transformer-based models, many researchers are keen to perform time series forecasting tasks using MLP-based models (Benidis et al. 2023). (Chen et al. 2023) proposed TSMixer with all-MLP architecture to efficiently utilize cross-variate and auxiliary information to improve the performance of time series forecasting. LightTS (Zhang et al. 2022) is dedicated to solving multivariate time series forecasting problems, and it can efficiently handle very long input series. (Yi et al. 2023) explores MLP in the frequency domain for time series forecasting and proposes a novel architecture for FreTS that includes two phases: domain conversion and frequency learning.
Compared to Transformer-based models, MLP-based models are simpler in structure, less complex and more efficient. However, the MLP-based models also suffer from a number of shortcomings. In the case of high volatility and non-periodic, non-stationary patterns, MLP performance relying only on past observed temporal patterns is not satisfactory (Chen et al. 2023). In other words, MLP-based models have low robustness. In addition, MLP is worse at capturing global dependencies compared to Transformers (Yi et al. 2023).
Applications of Mamba
Mamba (Gu and Dao 2023), as a new architecture, addresses to some extent the challenges faced by Transformer and MLP on time series forecasting tasks. It captures global dependencies better in a lightweight structure and has a better sense of position relationships. In addition, the Mamba architecture is more robust. Due to its excellent performance, Mamba swiftly attracted the attention of a large number of researchers. Pióro et al. (2024) replaced the Transformer architecture in the Mixture of Experts (MoE) with the Mamba architecture, achieving a complete override of Mamba’s and Transformer-MoE’s performance. Mamba has also been used to solve the long-range dependency problem in biomedical image segmentation tasks (Ma, Li, and Wang 2024). VideoMamba (Li et al. 2024) achieves efficient long-term modeling using Mamba’s linear complexity operator, showing advantages on long video understanding tasks. In addition, Mamba has demonstrated strong performance in clinical note generation (Yang et al. 2024), small target detection (Chen et al. 2024), etc.
However, to the best of our knowledge, there is still no work applying the Mamba architecture to time series forecasting tasks. Intuitively, the Mamba architecture effectively mitigates the problems faced by the Transformer and MLP architectures for time series forecasting. Therefore, we try to introduce the Mamba architecture in time series forecasting tasks.
Preliminaries
Problem Statement
In time series forecasting tasks, the model receives input as a history sequence $U_{in}=[u_{1},u_{2},\ldots,u_{L}]\in{\mathbb{R}^{L\times V}}$ and $u_{n}=[p_{1},p_{2},\ldots,p_{V}]$. and then uses this information to predict a future sequence $U_{out}=[u_{L+1},u_{L+2},\ldots,u_{L+T}]\in{\mathbb{R}^{T\times V}}$. The preceding $L$ and $T$ are referred to as the review window and prediction horizon respectively, representing the lengths of the past and future time windows, while $p$ is a variate and $V$ represents the total number of variates.
State Space Models
State space models can represent any cyclical process with latent states. By using first-order differential equations to represent the evolution of the system’s internal state and another set to describe the relationship between latent states and output sequences, input sequences $x(t)\in\mathbb{R}^{D}$ can be mapped to output sequences $y(t)\in\mathbb{R}^{N}$ through latent states $h(t)\in\mathbb{R}^{N}$ in (1):
$h(t)^{{}^{\prime}}$ $=\textbf{\emph{A}}h(t)+\textbf{\emph{B}}x(t),$ (1)
$y(t)$ $=\textbf{\emph{C}}h(t),$
[…]
$h_{t}$ $=\overline{\textbf{\emph{A}}}h_{t-1}+\overline{\textbf{\emph{B}}}x_{t},$ (2)
$y_{t}$ $=\textbf{\emph{C}}h_{t},$
where $\overline{\textbf{\emph{A}}}={\mathrm{exp}}(\Delta A)$ and $=(\Delta\textbf{\emph{A}})^{-1}({\mathrm{exp}}(\Delta\textbf{\emph{A}})-I)\cdot\Delta\textbf{\emph{B}}$ B. Since transitioning from continuous form $(\Delta,\textbf{\emph{A}},\textbf{\emph{B}},\textbf{\emph{C}})$ to discrete form $(\overline{\textbf{\emph{A}}},\overline{\textbf{\emph{B}}},\textbf{\emph{C}})$, the model can be efficiently calculated using a linear recursive approach (Gu et al. 2021). The structured state space model (S4) (Gu, Goel, and Ré 2021), originating from the vanilla SSM, utilizes HiPPO (Gu et al. 2020) for initialization to add structure to the state matrix A, thereby improving long-range dependency modeling.
Mamba (Gu and Dao 2023) introduces a data-dependent selection mechanism into the S4 and incorporates hardware-aware parallel algorithms in its looping mode, enabling Mamba to effectively capture contextual information in long sequences while maintaining computational efficiency. As an approximately linear perplexity series model, Mamba demonstrates significant potential in long sequence tasks, compared to transformers, in both efficiency enhancement and performance improvement.
Figure 1: Overall framework of S/D-Mamba, the left side of the figure presents the overall architecture of our model. The Mamba (sensitive) enclosed within the dotted line represents an additional Mamba block included in D-Mamba. The right side of the figure details the components of the S/D-Mamba Block.
Methodology
In this section, we provide a detailed introduction for S/D-Mamba. Fig.1 illustrates the overall structure of S/D-Mamba, which is primarily composed of three layers. The first layer, the embedding layer, utilizes a variable-based embedding approach. The second layer, the Mamba layer (S/D-Mamba block), forms the core of our model, capturing the dependencies among input variates’ embeddings and updating them according to variate correlations. The third layer, the Norm-FFN-Norm layer, further mixes the information output by the Mamba layer. Alg.1 demonstrates the operation process of S-Mamba. The procedure of D-Mamba remains unchanged, except for the addition of a parallel sensitive Mamba block within the Mamba Layer.
Algorithm 1 The Forecasting Procedure of S-Mamba
Input: $Batch(U_{in})=[u_{1},u_{2},\ldots u_{L}]:(B,L,V)$
Output: $Batch(U_{out})=[u_{L+1},u_{L+2},\ldots u_{L+T}]:(B,T,V)$
1:  Embedding Layer:
2:  $Batch(U_{in}^{\top}):(B,V,L)\leftarrow\mathrm{Transpose}(Batch(U_{in}))$Turn to variate side}
3:  $\bm{U}:(B,V,D)\leftarrow\mathrm{Linear}(Batch(U_{in}^{\top}))$Variate-based Embedding}
4:  for $l$ in $Layers$ do
5:     S-Mamba Layer:
6:     $x,z:(B,V,ED)\leftarrow\mathrm{Linear}(\bm{U})$Linear projection}
7:     $x^{{}^{\prime}}:(B,V,ED)\leftarrow\mathrm{SiLU}(\mathrm{Conv1D}(x))$
8:     $\textbf{\emph{A}}:(D,N)\leftarrow Parameter$Structured state matrix}
[…]
10:     $\Delta:(B,V,D)\leftarrow\mathrm{Softplus}(Parameter+\mathrm{Broadcast}(\mathrm{Linear}(x^{{}^{\prime}})))$
11:     $:(B.V.D.N)\leftarrow{discretize}(\Delta,\textbf{\emph{A}},\textbf{\emph{B}})$ A , B ) {Input-dependent parameters and discretization}
12:     $y:(B,V,ED)\leftarrow\mathrm{SelectiveSSM}(\overline{\textbf{\emph{A}}},\overline{\textbf{\emph{B}}},\textbf{\emph{C}})(x^{{}^{\prime}})$
13:     $y^{{}^{\prime}}:(B,V,ED)\leftarrow y\otimes\mathrm{SiLU}(z)$
14:     $\bm{Y}:(B,V,D)\leftarrow\mathrm{Linear}(y^{{}^{\prime}})$Linear Projection}
15:     Norm-FFN-Norm Layer:
16:     $\bm{U}^{{}^{\prime}}:(B,V,D)\leftarrow\mathrm{Dropout}(\bm{Y})+\bm{U}$Residual Connection}
17:     $\bm{U}^{{}^{\prime}}:(B,V,D)\leftarrow\mathrm{LayerNorm}(\bm{U}^{{}^{\prime}})$
18:     $\bm{U}^{{}^{\prime}}:(B,V,D)\leftarrow\mathrm{Feed-forward}(\bm{U}^{{}^{\prime}})$
19:     $\bm{U}^{{}^{\prime}}:(B,V,D)\leftarrow\mathrm{LayerNorm}(\bm{U}^{{}^{\prime}})$
20:  end for
21:  Projection:
22:  $\bm{U}^{{}^{\prime}}:(B,V,T)\leftarrow\mathrm{Projection}(\bm{U}^{{}^{\prime}})$
23:  $Batch(U_{out}):(B,T,V)\leftarrow\mathrm{Transpose}(\bm{U}^{{}^{\prime}})$
Embedding Layer
The input for the Embedding layer is $U_{in}$. Similar to iTransformer (Liu et al. 2023), we also set up a variate-based linear embedding layer, which considers each variate independently, thereby effectively reducing the noise caused by mixed variates according to time when embedding. After that, each variate information of the input data point is mapped into a high-dimensional space, where $D$ denotes the embedding dimension as (3).
$\bm{U}=\mathrm{Linear}(Batch(U_{in}))$ (3)
S/D-Mamba layer
The Mamba layer is considered the fundamental building block of the model architecture proposed in this paper. It functions to fuse information between variables based on VC.
Figure 2: The structure of selective SSM.
Mamba Block. The details are presented in the algorithm related to the mamba layer in Alg.1 and the description in Fig. 2. The Mamba layer takes a sequence $\bm{U}\in\mathbb{R}^{B\times V\times D}$ as input at time step $t$, where $B$ denotes the batch size, $V$ denotes the number of variates, and $D$ denotes hidden dimension. The block first expands the hidden dimension to $ED$ through linear projection, obtaining $x$ and $z$. Then, it processes the projection obtained earlier using convolutional functions and a SiLU (Elfwing, Uchibe, and Doya 2017) activation function. Based on the discretized SSM selected by the input parameters, denoted as the core of the mamba block, together with $x^{\prime}$, it generates the state representation $y$. Finally, $y$ is combined with a residual connection from $z$ after activation, and the final output $y_{t}$ at time step $t$ is obtained through a linear transformation. In summary, the Mamba Block effectively handles sequential information by leveraging selective state space models and input-dependent adaptations. The parameters involved in the Mamba Block include an SSM state expansion factor $N$, a size of convolutional kernel $k$, and a block expansion factor $E$ for input-output linear projection. The larger the values of $N$ and $E$, the higher the computational cost. The final output of the Mamba block is $\bm{Y}\in\mathbb{R}^{B\times V\times D}$. When there are two Mamba Blocks (D-Mamba), the inputs of the two mamba blocks are aggregated as the output $\bm{Y}=\bm{Y_{1}}+\bm{Y_{2}}$, where $\bm{Y_{1}}$ and $\bm{Y_{2}}$ are the outputs of the two mamba layers respectively.
Norm-FFN-Norm Layer
At this layer, we further process the output of the Mamba Layer. Firstly, we employ a normalization layer (Liu et al. 2023), originally designed to enhance convergence and training stability in deep networks. Prior to layer normalization, we standardize all variables to a Gaussian distribution, thereby minimizing disparities resulting from inconsistent measurements. Then, the feed-forward network (FFN) is used on the series representation of each variate, allowing for complex time series descriptions. FFN encodes observed time series and decodes future series representations using dense non-linear connections, similar to recent MLP-based approaches. Finally, another normalization layer same as the former is set to adjust the future series representations.
Table 1: Full results of the long-term time series forecasting task. We compare the representative and the state-of-the-art models with S-Mamba and D-Mamba under different prediction lengths following the setting of iTransformer. The lookback length $L$ is set to 96 for all models. The number on the left side of the table represents the length of the output sequence $T$. The optimal results are highlighted in bold red font, while the suboptimal results are presented in underlined blue font. The results of baselines are reported by Liu et al. (2023).
Models D-Mamba S-Mamba iTransformer RLinear PatchTST DLinear FEDformer Autoformer
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
PEMS03 12 0.066 0.171 0.066 0.171 0.071 0.174 0.126 0.236 0.099 0.216 0.122 0.243 0.126 0.251 0.272 0.385
24 0.087 0.196 0.088 0.197 0.093 0.201 0.246 0.334 0.142 0.259 0.201 0.317 0.149 0.275 0.334 0.440
48 0.133 0.243 0.138 0.248 0.125 0.236 0.551 0.529 0.211 0.319 0.333 0.425 0.227 0.348 1.032 0.782
96 0.197 0.301 0.217 0.322 0.164 0.275 1.057 0.787 0.269 0.370 0.457 0.515 0.348 0.434 1.031 0.796
PEMS04 12 0.073 0.178 0.075 0.179 0.078 0.183 0.138 0.252 0.105 0.224 0.148 0.272 0.138 0.262 0.424 0.491
24 0.085 0.193 0.087 0.196 0.095 0.205 0.258 0.348 0.153 0.275 0.224 0.340 0.177 0.293 0.459 0.509
48 0.104 0.214 0.105 0.217 0.120 0.233 0.572 0.544 0.229 0.339 0.355 0.437 0.270 0.368 0.646 0.610
96 0.125 0.234 0.128 0.239 0.150 0.262 1.137 0.820 0.291 0.389 0.452 0.504 0.341 0.427 0.912 0.748
PEMS07 12 0.062 0.160 0.064 0.162 0.067 0.165 0.118 0.235 0.095 0.207 0.115 0.242 0.109 0.225 0.199 0.336
24 0.079 0.183 0.081 0.184 0.088 0.190 0.242 0.341 0.150 0.262 0.210 0.329 0.125 0.244 0.323 0.420
48 0.091 0.188 0.094 0.193 0.110 0.215 0.562 0.541 0.253 0.340 0.398 0.458 0.165 0.288 0.390 0.470
96 0.197 0.301 0.199 0.305 0.139 0.245 1.096 0.795 0.346 0.404 0.594 0.553 0.262 0.376 0.554 0.578
PEMS08 12 0.084 0.189 0.085 0.189 0.079 0.182 0.133 0.247 0.168 0.232 0.154 0.276 0.173 0.273 0.436 0.485
24 0.126 0.232 0.127 0.233 0.115 0.219 0.249 0.343 0.224 0.281 0.248 0.353 0.210 0.301 0.467 0.502
48 0.180 0.228 0.196 0.229 0.186 0.235 0.569 0.544 0.321 0.354 0.440 0.470 0.320 0.394 0.966 0.733
96 0.243 0.254 0.245 0.255 0.221 0.267 1.166 0.814 0.408 0.417 0.674 0.565 0.442 0.465 1.385 0.915
ETTm1 96 0.329 0.364 0.325 0.361 0.334 0.368 0.355 0.376 0.329 0.367 0.345 0.372 0.379 0.419 0.505 0.475
192 0.371 0.387 0.368 0.385 0.377 0.391 0.391 0.392 0.367 0.385 0.380 0.389 0.426 0.441 0.553 0.496
336 0.403 0.409 0.401 0.408 0.426 0.420 0.424 0.415 0.399 0.410 0.413 0.413 0.445 0.459 0.621 0.537
720 0.471 0.448 0.469 0.446 0.491 0.459 0.487 0.450 0.454 0.439 0.474 0.453 0.543 0.490 0.671 0.561
ETTm2 96 0.179 0.262 0.180 0.264 0.180 0.264 0.182 0.265 0.175 0.259 0.193 0.292 0.203 0.287 0.255 0.339
192 0.245 0.307 0.245 0.306 0.250 0.309 0.246 0.304 0.241 0.302 0.284 0.362 0.269 0.328 0.281 0.340
336 0.309 0.347 0.309 0.347 0.311 0.348 0.307 0.342 0.305 0.343 0.369 0.427 0.325 0.366 0.339 0.372
720 0.406 0.402 0.408 0.403 0.412 0.407 0.407 0.398 0.402 0.400 0.554 0.522 0.421 0.415 0.433 0.432
ETTh1 96 0.385 0.405 0.382 0.401 0.386 0.405 0.386 0.395 0.414 0.419 0.386 0.400 0.376 0.419 0.449 0.459
192 0.433 0.433 0.434 0.435 0.441 0.436 0.437 0.424 0.460 0.445 0.437 0.432 0.420 0.448 0.500 0.482
336 0.475 0.455 0.482 0.458 0.487 0.458 0.479 0.446 0.501 0.466 0.481 0.459 0.459 0.465 0.521 0.496
720 0.482 0.479 0.481 0.480 0.503 0.491 0.481 0.470 0.500 0.488 0.519 0.516 0.506 0.507 0.514 0.512
ETTh2 96 0.294 0.348 0.295 0.348 0.297 0.349 0.288 0.338 0.302 0.348 0.333 0.387 0.358 0.397 0.346 0.388
192 0.376 0.398 0.376 0.397 0.380 0.400 0.374 0.390 0.388 0.400 0.477 0.476 0.429 0.439 0.456 0.452
336 0.417 0.430 0.420 0.432 0.428 0.432 0.415 0.426 0.426 0.433 0.594 0.541 0.496 0.487 0.482 0.486
720 0.434 0.449 0.425 0.441 0.427 0.445 0.420 0.440 0.431 0.446 0.831 0.657 0.463 0.474 0.515 0.511
Electricity 96 0.142 0.238 0.144 0.239 0.148 0.240 0.201 0.281 0.181 0.270 0.197 0.282 0.193 0.308 0.201 0.317
192 0.160 0.253 0.160 0.252 0.162 0.253 0.201 0.283 0.188 0.274 0.196 0.285 0.201 0.315 0.222 0.334
336 0.176 0.271 0.176 0.271 0.178 0.269 0.215 0.298 0.204 0.293 0.209 0.301 0.214 0.329 0.231 0.338
720 0.203 0.297 0.207 0.300 0.225 0.317 0.257 0.331 0.246 0.324 0.245 0.333 0.246 0.355 0.254 0.361
Exchange 96 0.088 0.209 0.087 0.209 0.086 0.206 0.093 0.217 0.088 0.205 0.088 0.218 0.148 0.278 0.197 0.323
192 0.184 0.306 0.182 0.304 0.177 0.299 0.184 0.307 0.176 0.299 0.176 0.315 0.271 0.315 0.300 0.369
336 0.334 0.419 0.336 0.421 0.331 0.417 0.351 0.432 0.301 0.397 0.313 0.427 0.460 0.427 0.509 0.524
720 0.866 0.702 0.857 0.698 0.847 0.691 0.886 0.714 0.901 0.714 0.839 0.695 1.195 0.695 1.447 0.941
Traffic 96 0.387 0.260 0.387 0.262 0.395 0.268 0.649 0.389 0.462 0.295 0.650 0.396 0.587 0.366 0.613 0.388
192 0.408 0.269 0.410 0.270 0.417 0.276 0.601 0.366 0.466 0.296 0.598 0.370 0.604 0.373 0.616 0.382
336 0.423 0.276 0.425 0.277 0.433 0.283 0.609 0.369 0.482 0.304 0.605 0.373 0.621 0.383 0.622 0.337
720 0.458 0.294 0.459 0.296 0.467 0.302 0.647 0.387 0.514 0.322 0.645 0.394 0.626 0.382 0.660 0.408
Weather 96 0.170 0.212 0.172 0.214 0.174 0.214 0.192 0.232 0.177 0.218 0.196 0.255 0.217 0.296 0.266 0.336
192 0.220 0.257 0.221 0.259 0.221 0.254 0.240 0.271 0.225 0.259 0.237 0.296 0.276 0.336 0.307 0.367
336 0.276 0.298 0.280 0.300 0.278 0.296 0.292 0.307 0.278 0.297 0.283 0.335 0.339 0.380 0.359 0.395
720 0.356 0.348 0.358 0.349 0.358 0.347 0.364 0.353 0.354 0.348 0.345 0.381 0.403 0.428 0.419 0.428
Solar 96 0.202 0.233 0.208 0.241 0.203 0.237 0.322 0.339 0.234 0.286 0.290 0.378 0.242 0.342 0.884 0.711
192 0.246 0.263 0.232 0.259 0.233 0.261 0.359 0.356 0.267 0.310 0.320 0.398 0.285 0.380 0.834 0.692
336 0.257 0.263 0.262 0.281 0.248 0.273 0.397 0.369 0.290 0.315 0.353 0.415 0.282 0.376 0.941 0.723
720 0.256 0.283 0.258 0.284 0.249 0.275 0.397 0.356 0.289 0.317 0.356 0.413 0.357 0.427 0.882 0.717
$1^{\text{st}}$ Count 22 19 7 13 9 6 5 10 9 7 3 0 3 0 0 0
$2^{\text{nd}}$ Count 18 19 23 23 9 8 1 1 3 5 1 3 0 1 0 0
Experiments
Datasets and Baselines
We conduct extensive experiments on several real-world datasets, including PEMS (PEMS03, PEMS04, PEMS07, PEMS08), ETT (Electricity Transformer Temperature) (Zhou et al. 2021) (ETTh1, ETTh2, ETTm1, ETTm2), Electricity, Exchange-Rate (Lai et al. 2018), Traffic, Weather, Solar-Energy. All of them are multivariate time series datasets adopted by Liu et al. (2023). Tab. 2 shows the statistics of these datasets.
Our models are fairly compared with 6 representative and state-of-the-art forecasting models, including (1) Transformer-based methods: Autoformer (Wu et al. 2021), FEDformer (Zhou et al. 2022), PatchTST (Nie et al. 2022), and (2) Linear-based methods: DLinear (Zeng et al. 2023), RLinear (Li et al. 2023b).
Table 2: The statistics of the datasets.
Datasets Variates Timesteps Granularity
PEMS03 358 26,209 5min
PEMS04 307 16,992 5min
PEMS07 883 28,224 5min
PEMS08 170 17,856 5min
ETTm1 &\&& ETTm2 7 17,420 15min
ETTh1 &\&& ETTh2 7 69,680 1hour
Electricity 321 26,304 1hour
Exchange-Rate 8 7,588 1day
Traffic 862 17,544 1hour
Weather 21 52,696 10min
Solar-Energy 137 52,560 10min
Overall Performance
Tab.1 presents a comparative analysis of the overall performance of our models and other baseline models across all datasets. From the data presented in the table, we can deduce the following three observations: (1) The D-Mamba and S-Mamba models generally maintain a performance advantage, albeit a slight one, in most scenarios. (2) D-Mamba outperforms S-Mamba in the majority of cases, suggesting that in TSF tasks, it may be more effective to configure Mamba with varying sensitivities to work collaboratively when extracting relationships between variables. (3) The distinction between S-Mamba and iTransformer lies in their respective approaches to extracting dependencies between variates. On certain datasets, substituting Transformer with Mamba leads to a decrease in effectiveness. This raises an intriguing research question: under what circumstances does Mamba surpass Transformer in performance?
Table 3: Comparison of costs between S/D-Mamba and three baselines, including GPU memory usage, and training time on ECL and Traffic. The number on the left side of the table represents the lookback length $L\in\{96,192,336,720\}$, while the forecasting length $T=96$.
Models D-Mamba S-Mamba iTransformer RLinear Autoformer
Metric Memory Speed Memory Speed Memory Speed Memory Speed Memory Speed
(GiB) (ms/iters) (GiB) (ms/iters) (GiB) (ms/iters) (GiB) (ms/iters) (GiB) (ms/iters)
ECL 96 0.537 39.5 0.402 37.2 0.771 41.5 0.042 18.1 1.696 81.6
192 0.540 40.3 0.411 38.0 0.775 43.6 0.054 22.8 2.426 105.0
336 0.549 41.3 0.427 38.6 0.782 45.5 0.071 30.6 3.702 148.4
720 0.564 43.4 0.467 39.4 0.800 56.8 0.115 47.0 6.991 247.2
Traffic 96 1.667 11.3 1.262 97.3 4.549 161.2 0.085 39.3 1.791 203.2
192 1.676 11.5 1.288 97.5 4.558 154.6 0.115 50.7 2.532 215.2
336 1.725 11.8 1.326 101.5 4.575 160.9 0.161 76.5 3.818 277.9
720 1.749 11.9 1.427 116.6 4.617 167.4 0.279 137.4 7.133 398.3
Model Efficiency
To evaluate the computational efficiency of the models, We compare the memory usage and computing time of S/D Mamba with several baselines on the ECL and Traffic. Specifically, we conduct independent runs on a single NVIDIA A6000 GPU and meticulously document the results in Tab.3. The numbers in the left column of the table represent the lookback length $L\in\{96,192,336,720\}$ The results indicate that, in comparison to the Transformer-based model, both S-Mamba and D-Mamba exhibit optimization in terms of time cost and GPU memory usage, and demonstrate a performance more akin to that of linear models, R-Linear.
Table 4: Performance comparison between Autoformer, Flashformer and Flowformer and their variants with Mamba encoder block on ECL and Traffic datasets, where we fixed the lookback length $L=96$ and forcasting length $T=96$.
Datasets ECL Traffic
Models Auto Auto-M Flash Flash-M Flow Flow-M Auto Auto-M Flash Flash-M Flow Flow-M
Memory(GiB) 1.59 1.22 0.91 0.63 0.83 0.68 1.78 1.35 1.16 0.80 1.06 0.79
Speed(ms/iters) 76.3 56.1 62.8 40.6 57.4 38.0 126.1 97.3 110.3 91.5 101.5 81.2
MSE 0.201 0.191 0.253 0.252 0.252 0.246 0.613 0.607 0.68 0.67 0.658 0.649
MAE 0.317 0.305 0.347 0.345 0.349 0.344 0.338 0.381 0.374 0.465 0.360 0.365
Is Mamba a better Variate-Encoder?
In prior studies, the encoder of the Transformer is commonly utilized for variant-encoder tasks, which involve aggregating all variable information at a single time step. From a structural perspective, in most Transformer-based models with encoder-decoder architectures, the self-attention mechanism within the encoder layer can be substituted with the Mamba mechanism. Furthermore, from a utility perspective, Mamba’s selective mechanism can distinguish VC among different variables, enabling it to function as a Variate-Encoder. Therefore, we set an experiment by replacing the encoder layer (self-attention) of Transformer (Vaswani et al. 2017), Autoformer (Wu et al. 2021), Flashformer (Dao et al. 2022) and Flowformer (Wu et al. 2022) with a simple Mamba block to get Trans-M, Auto-M, Fed-M, Flash-M and Flow-M and and compare their performance. The results are astonishing as 4, which straightforward modification with Mamba enhances the performance of these models and also reduces GPU memory usage and running time, which may indicate that Mamba has better results when aggregating multi-variates information.
Can Mamba help learn temporal information?
Prior research has shown that the performance of Transformer-based models does not consistently improve with increasing lookback sequence length $L$, which is somewhat unexpected. A plausible explanation is that the temporal sequence relationship is overlooked under the self-attention mechanism, as it disregards the sequential order, and in some instances, even inverts it. Mamba, resembling a Recurrent Neural Network (Medsker, Jain et al. 2001), concentrates on the preceding window during information extraction, thereby preserving certain sequential attributes. This prompts an exploration of Mamba’s potential effectiveness in temporal sequence information fusion, aiming to address the issue of diminishing or stagnant performance with increasing lookback length. Consequently, we add an additional Mamba block between the encoder Layer and decoder layer of Transformers. The role of this Mamba Block is to add a layer of time sequence dependence from the information output by the encoder layer, to add some information similar to position embedding before the decoder layer processes it. We experiment with Autoformer (Wu et al. 2021), Transformer (Vaswani et al. 2017), and Informer (Zhou et al. 2021) to get Auto-M, Infor-M, and Refor-M, and evaluate their performance with varying lookback lengths. We also test the performance of S/D-Mamba as the lookback length changes. The results are in the Fig.3, from which we can observe three results. (1) S/D-Mamba can enhance its performance as the input lengthens, but we believe this is not solely due to the Mamba Block, but rather to the Variate-based Embedding method proposed by iTransformer (Liu et al. 2023). (2) With the addition of a Mamba Block, all three models show performance improvements. Notably, the performance of Informer has seen the most significant enhancement. (3) Despite these variants’ performance gains sometimes, they do not achieve optimization with longer lookback lengths. This is consistent with the findings of Zeng et al. (2023), which also suggest that encoding temporal sequence information into the model beforehand does not resolve this issue.
Figure 3: Forecasting performance on ECL and Traffic with the lookback length $L\in\{96,192,336,720\}$ while the prediction length $S=96$.
Is Mamba generalizable in TSF?
The emergence of pre-trained models (Devlin et al. 2018) and large language models (Chang et al. 2023) based on the Transformer architecture has underscored the Transformer’s ability to discern similar patterns across diverse data, highlighting its generalization capabilities. In the context of TSF, it is observed that all variates exhibit a similar pattern of differences, for example, temperature and humidity in weather forecasts always have similar trends. In other words, the generalization potential of the Transformer for sequence data may also take effect on TSF tasks.
Figure 4: Forecasting performance comparison between S/D-Mamba and iTransformer trained on 100% variates with on 40% variates. The lookback length $L=96$ for all datasets. For the PEMS dataset, the output length $T=12$, while for the other datasets $T=96$.
In this vein, iTransformer (Liu et al. 2023) conducts a pivotal experiment. The study involves masking a majority of the variates in a dataset and training the model on a limited subset of variates. Subsequently, the model was tasked with forecasting all variates, including those previously unseen, based on the learned information from the few variates. Building on this, we seek to evaluate the generalization capabilities of Mamba in TSF tasks. An experiment is proposed wherein the S/D-Mamba are trained on merely 40% of the variates in the PEMS03, PEMS04, ECL, Weather, Traffic, and Solar datasets. Then they are employed to predict 100% variates, and the results are subjected to statistical analysis. The outcomes of this investigation in 4reveal that both S-Mamba and D-Mamba exhibit generalization potential in the six datasets, which proves their generalizability in TSF tasks. However, under this training paradigm, they demonstrate inferior performance compared to iTransformer, underscoring their generalizability in TSF tasks.
Conlusion
Transformer-based models have consistently exhibited outstanding performance in the field of time series forecasting (TSF), while Mamba has recently gained popularity, and has been shown to surpass the Transformer in various domains by delivering superior performance while reducing memory and computational overhead. Motivated by these advancements, we sought to investigate the potential of Mamba-based models in the TSF domain, with the aim of uncovering new research avenues for this field. To this end, we introduced two variants of the Mamba model, S-Mamba and D-Mamba, which transfer the task of variable information fusion from the Transformer architecture to the Mamba block. We evaluated their performance across a multitude of datasets, and our findings indicate that S-Mamba and D-Mamba not only reduce GPU memory and computational overhead but also achieve leading performance in the TSF domain. Furthermore, we conduct extensive experiments to explore the extent to which Mamba can outperform the Transformer. The results prove Mamba possesses robust capabilities and exhibits remarkable potential to replace Transformer in the TSF tasks.